A carriage of mass 150 kg moving with a velocity of 1.2 m/s collides with a 300 kg carriage that is stationary on track. The second carriage moves at 0.8 m/s after the collision. Calculate the velocity of the first carriage after the collision.
1
Expert's answer
2020-03-24T09:57:16-0400

From the conservation of momentum:


"mv=Mu+mv'"

"(150)(1.2)=(300)(0.8)+150v'"

"v'=-0.4\frac{m}{s}"